# Changing the World?

### v1

@@ 2.2.1 Literature Search and Curation
Building on this spectrum of approaches, researchers have developed varied implementations for searching and filtering scientific literature. Schmidgall et al. (2025) utilize the arXiv API to retrieve papers and perform summarization, full-text extraction, and curation tasks. Meanwhile, Jiabin et al. (2025) apply citation-based filtering to top-cited domain papers from arXiv. Beyond single-agent retrieval, multi-agent systems have been explored as well. Gottweis et al. (2025) integrate web search tools, while Lu et al. (2024) utilize Semantic Scholar’s API to identify relevant papers for hypotheses generation. In the RAG domain, Lála et al. (2023) and Skarlinski et al. (2024) introduce PaperQA and PaperQA2 for scientific questioning answering tasks. Furthermore, the pursuit of greater autonomy is exemplified by PaSa (He et al., 2025), which deploys a dual-agent architecture with session-level reinforcement learning to autonomously search, analyze papers, and explore citation networks for complex queries. Despite these advancements, Beel et al. (2025) note that existing systems still heavily rely on keyword-based matching, often retrieving classic but potentially outdated papers, highlighting the need for more context-aware retrieval strategies.
@@ 7.2 Limitations of Current AI Scientists in Scientific Discovery
Despite the substantial progress in the field of AI Scientist systems, there is still a considerable gap before AI Scientist can mature into scientific agents capable of making ground-breaking discoveries, thus changing the world and reshaping the scientific research paradigm. This persistent gap can be attributed to several critical factors. First, the inherent limitations of the foundation models (i.e., LLMs) that underpin AI Scientist systems pose significant barriers. These models, while demonstrating impressive capabilities in language understanding and generation, often lack the depth of domain expertise, capacity for scientific reasoning, and nuanced judgment required for high-impact scientific breakthroughs. Their outputs are constrained by the knowledge encoded in their training data, which may be outdated, incomplete, or not fully aligned with the latest advancements in the scientific community. Second, current AI Scientist systems exhibit inadequate scientific research abilities in several key aspects. For instance, they may generate hypotheses that lack feasibility or novelty, apply inappropriate or insufficient validation methods, and struggle with the deep comprehension and decomposition of complex research problems. These shortcomings not only limit the reliability of AI-generated scientific artifacts but also hinder the AI Scientist systems’ ability to independently advance the frontiers of science.
@@ 7.2.1 Fundamental Limitations of Base Models
High cost and inefficiency of knowledge updating. Cutting-edge scientific research advances rapidly, requiring AI Scientist systems to constantly incorporate the latest findings and data constantly. However, keeping LLMs up-to-date is non-trivial: (1) The volume of new scientific literature and data is immense, and integrating this information into an LLM demands significant computational resources, storage, and time. (2) Frequent retraining or fine-tuning of large-scale models is both costly and technically challenging, often leading to delays in knowledge integration that may render the model outdated in fast-moving research domains. (3) The lack of efficient, fine-grained, and continuous learning mechanisms means that knowledge refreshing processes are typically slow, creating a persistent lag between the state of the art in scientific fields and the knowledge encoded in AI Scientist systems.
Catastrophic Forgetting. Another fundamental challenge is catastrophic forgetting, which is the tendency of neural networks to lose previously acquired information when trained on new data (Kirkpatrick et al., 2017). For AI Scientist systems, this challenge means that attempts to update the model with new scientific knowledge may inadvertently overwrite or degrade its understanding of existing scientific concepts, methods, and facts. Such instability undermines the system’s ability to maintain a comprehensive and coherent body of scientific knowledge over time. As a result, integrating findings across different periods and scientific fields becomes particularly challenging, which in turn impairs the LLM-based AI Scientist systems’ capacities to perform longitudinal analyses, connect historical insights with emerging discoveries, and deliver consistent, reliable support for complex scientific explorations.
@@ 7.2.2 Limitations of Research Capabilities of AI Scientist Systems
Knowledge acquisition. The ability to acquire, organize, and internalize knowledge from the vast and rapidly expanding body of scientific literature is a fundamental capability for an AI Scientist. This includes not only searching for and retrieving relevant papers and results, but also extracting core ideas, incorporating experimental details, and integrating disparate findings into a coherent understanding. However, existing AI Scientist systems face several substantial challenges in this area, including: (1) Precision in information retrieval. The enormous scale and diversity of scientific literature make it difficult for current AI Scientist systems to accurately retrieve the most relevant and up-to-date information for a given research problem (Landhuis, 2016). They often struggle with the precise interpretation of complex or ambiguous queries, effective filtering of highly relevant documents, and ensuring that the results remain timely and reliable. Both precision and recall in retrieval must be further improved to meet scientific standards. (2) Multi-document summarization and synthesis. Generating accurate and comprehensive literature reviews requires not only retrieving multiple relevant documents, but also synthesizing information across sources, resolving inconsistencies, and generating faithful summaries (Agarwal et al., 2024a, b). Recent approaches, such as RAG and plan-based summarization, have made progress in this area. However, studies have shown that LLMs are still prone to misquoting or distorting source materials (Agarwal et al., 2024b), indicating persistent challenges in semantic understanding, information integration, and factual accuracy when summarizing scientific knowledge from multiple documents.
@@
Idea generation. One of the most critical aspects of scientific research is the generation of novel and feasible scientific hypotheses. Despite notable advancements, the idea generation capability remains significantly constrained by several limitations: (1) Difficulty in generating high-quality hypotheses. Generating scientific hypotheses that are not only novel but also feasible is a complex task that requires a deep understanding of both existing knowledge and the ability to identify innovative connections. However, empirical evidence suggests that current AI Scientist systems often fall short in this regard. Specifically, the ideas produced by these models tend to lack true originality and are frequently repetitive across different runs or even across different models (Lu et al., 2024). The underlying cause of this phenomenon is twofold: First, LLMs are fundamentally limited by their training data, which constrains their ability to move beyond well-trodden conceptual ground. Second, without advanced mechanisms for knowledge integration or creativity, AI Scientist systems tend to rely heavily on superficial text correlations, rather than generating ideas with genuine scientific novelty. This not only leads to repetitive outputs but also limits the potential for breakthrough discoveries. (2) Challenges in evaluating the quality of AI-generated scientific hypotheses. A further limitation arises from the challenge of reliably assessing the quality and potential impact of AI-generated scientific ideas. Unlike standardized tasks with clear evaluation criteria, the value of a scientific hypothesis is often subjective, context-dependent, and difficult to quantify. Current evaluation methods rely heavily on human expert judgment, which is not only resource-intensive but also susceptible to subjectivity and inconsistency. Recent attempts to introduce automated evaluators have helped to some extent, but these systems themselves inherit the biases and knowledge limitations of their underlying models. As a result, the lack of reliable, objective, and scalable evaluation frameworks for AI-generated hypotheses remains a bottleneck for the advancement of AI Scientist systems in real-world research settings.
@@ 8.2.1 AI Scientist for Individual Researcher
Preference alignment of human researchers. Effective alignment involves matching the capabilities of AI Scientist systems with human research goals, including but not limited to: (1) Value-driven hypothesis generation: aligning hypothesis proposals with the researcher’s research preferences, risk tolerance and resource constraints. (2) Ethical boundary management: Dynamically adapt to the researcher’s institutional rules, funding ethics, and personal boundaries (e.g., auto-rejecting high-risk topics like gene editing). (3) Collaborative research agendas: jointly planning research directions with human researchers, taking into account individual career aspirations and priorities for impactful scientific contributions.
@@ 8.2.2 AI Scientist for Human Society
Prospects for human development. AI Scientist systems can greatly boost human progress by: k (Romera-Paredes et al., 2024) Reducing resource gaps by providing open-access tools and cloud-based experimental simulators, enabling under-resourced institutions to participate in global research. k (Romera-Paredes et al., 2024) Accelerating discovery for critical challenges (e.g., renewable energy, disease treatment) through automated experimentation. Early successes like ChemCrow (Bran et al., 2023) and BioDiscoveryAgent (Roohani et al., 2024) show promise in chemistry/biology. As these systems increasingly interface with physical laboratories (e.g., IoT sensors), they may evolve to support full-cycle scientific workflows.
@@ 9 Conclusion
This paper presented a systematic survey of existing research on AI Scientist systems, providing a comprehensive overview of advances in the field. Specifically, we proposed a capability-level framework that divides the capabilities of an AI Scientist into four progressive stages: knowledge acquisition, idea generation, verification and falsification, and evolution. Based on this framework, we critically analyzed current research and highlighted gaps AI Scientist systems must address before they can make ground-breaking discoveries that solve grand challenges across medicine, energy, and the environment, thereby changing the world and reshaping the scientific research paradigm. Finally, we concluded the survey by outlining future directions essential for closing the gap and advancing the field toward truly autonomous scientific intelligence.

### v2

@@ 2.2.1 Literature Search and Curation
Building on this spectrum of approaches, researchers have developed varied implementations for searching and filtering scientific literature. Schmidgall et al. (2025) utilize the arXiv API to retrieve papers and perform summarization, full-text extraction, and curation tasks. Meanwhile, Jiabin et al. (2025) apply citation-based filtering to top-cited domain papers from arXiv. Beyond single-agent retrieval, multi-agent systems have been explored as well. Gottweis et al. (2025) integrate web search tools, while Lu et al. (2024) utilize Semantic Scholar’s API to identify relevant papers for hypotheses generation. In the RAG domain, Lála et al. (2023) and Skarlinski et al. (2024) introduce PaperQA and PaperQA2 for scientific question answering tasks. Furthermore, the pursuit of greater autonomy is exemplified by PaSa (He et al., 2025), which deploys a dual-agent architecture with session-level reinforcement learning to autonomously search, analyze papers, and explore citation networks for complex queries. Despite these advancements, Beel et al. (2025) note that existing systems still heavily rely on keyword-based matching, often retrieving classic but potentially outdated papers, highlighting the need for more context-aware retrieval strategies.
@@ 7.2 Limitations of Current AI Scientists in Scientific Discovery
Despite the substantial progress in the field of AI Scientist systems, there is still a considerable gap before AI Scientist can mature into scientific agents capable of making ground-breaking discoveries, thus changing the world and reshaping the scientific research paradigm. This persistent gap can be attributed to several critical factors. First, the inherent limitations of the foundation models (i.e., LLMs) that underpin AI Scientist systems pose significant barriers. These models, while demonstrating impressive capabilities in language understanding and generation, often lack the depth of domain expertise, capacity for scientific reasoning, and nuanced judgment required for high-impact scientific breakthroughs. Their outputs are constrained by the knowledge encoded in their training data, which may be outdated, incomplete, or not fully aligned with the latest advancements in the scientific community. Second, current AI Scientist systems exhibit inadequate scientific research abilities in several key aspects. For instance, they may generate hypotheses that lack feasibility or novelty, apply inappropriate or insufficient validation methods, and struggle with the deep comprehension and decomposition of complex research problems. These shortcomings not only limit the reliability of AI-generated scientific artifacts but also hinder the AI Scientist systems’ abilities to independently advance the frontiers of science.
@@ 7.2.1 Fundamental Limitations of Base Models
High cost and inefficiency of knowledge updating. Cutting-edge scientific research advances rapidly, requiring AI Scientist systems to constantly incorporate the latest findings and data. However, keeping LLMs up-to-date is non-trivial: (1) The volume of new scientific literature and data is immense, and integrating this information into an LLM demands significant computational resources, storage, and time. (2) Frequent retraining or fine-tuning of large-scale models is both costly and technically challenging, often leading to delays in knowledge integration that may render the model outdated in fast-moving research domains. (3) The lack of efficient, fine-grained, and continuous learning mechanisms means that knowledge refreshing processes are typically slow, creating a persistent lag between the state of the art in scientific fields and the knowledge encoded in AI Scientist systems.
Catastrophic Forgetting. Another fundamental challenge is catastrophic forgetting, which is the tendency of neural networks to lose previously acquired information when trained on new data (Kirkpatrick et al., 2017). For an AI Scientist, this challenge means that attempts to update the model with new scientific knowledge may inadvertently overwrite or degrade its understanding of existing scientific concepts, methods, and facts. Such instability undermines the system’s ability to maintain a comprehensive and coherent body of scientific knowledge over time. As a result, integrating findings across different periods and scientific fields becomes particularly challenging, which in turn impairs the LLM-based AI Scientist systems’ capacities to perform longitudinal analyses, connect historical insights with emerging discoveries, and deliver consistent, reliable support for complex scientific explorations.
@@ 7.2.2 Limitations of Research Capabilities of AI Scientist Systems
Knowledge acquisition. The ability to acquire, organize, and internalize knowledge from the vast and rapidly expanding body of scientific literature is the fundamental capability for an AI Scientist. This includes not only searching for and retrieving relevant papers and results, but also extracting core ideas, incorporating experimental details, and integrating disparate findings into a coherent understanding. However, existing AI Scientist systems face several substantial challenges in this area, including: (1) Precision in information retrieval. The enormous scale and diversity of scientific literature make it difficult for current AI Scientist systems to accurately retrieve the most relevant and up-to-date information for a given research problem (Landhuis, 2016). They often struggle with the precise interpretation of complex or ambiguous queries, effective filtering of highly relevant documents, and ensuring that the results remain timely and reliable. Both precision and recall in retrieval must be further improved to meet scientific standards. (2) Multi-document summarization and synthesis. Generating accurate and comprehensive literature reviews requires not only retrieving multiple relevant documents, but also synthesizing information across sources, resolving inconsistencies, and generating faithful summaries (Agarwal et al., 2024a). Recent approaches, such as RAG and plan-based summarization, have made progress in this area. However, studies have shown that LLMs are still prone to misquoting or distorting source materials (Agarwal et al., 2024b), indicating persistent challenges in semantic understanding, information integration, and factual accuracy when summarizing scientific knowledge from multiple documents.
@@
Idea generation. One of the most critical aspects of scientific research is the generation of novel and feasible scientific hypotheses. Despite notable advancements, the idea generation capability remains significantly constrained by several limitations: (1) Difficulty in generating high-quality hypotheses. Generating scientific hypotheses that are not only novel but also feasible is a complex task that requires a deep understanding of both existing knowledge and the ability to identify innovative connections. However, empirical evidence suggests that current AI Scientist systems often fall short in this regard. Specifically, the ideas produced by these models tend to lack true originality and are frequently repetitive across different runs or even across different models (Lu et al., 2024). The underlying cause of this phenomenon is twofold: First, LLMs are fundamentally limited by their training data, which constrains their ability to move beyond well-trodden conceptual ground. Second, without advanced mechanisms for knowledge integration or creativity, AI Scientist systems tend to rely heavily on superficial text correlations, rather than generating ideas with genuine scientific novelty. This not only leads to repetitive outputs but also limits the potential for breakthrough discoveries. (2) Challenges in evaluating the quality of AI-generated scientific hypotheses. A further limitation arises from the challenge of reliably assessing the quality and potential impact of AI-generated scientific hypotheses. Unlike standardized tasks with clear evaluation criteria, the value of a scientific hypothesis is often subjective, context-dependent, and difficult to quantify. Current evaluation methods rely heavily on human expert judgment, which is not only resource-intensive but also susceptible to subjectivity and inconsistency. Recent attempts to introduce automated evaluators have helped to some extent, but these systems themselves inherit the biases and knowledge limitations of their underlying models. As a result, the lack of reliable, objective, and scalable evaluation frameworks for AI-generated hypotheses remains a bottleneck for the advancement of AI Scientist systems in real-world research settings.
@@ 8.2.1 AI Scientist for Individual Researcher
Preference alignment of human researchers. Effective alignment involves matching the capabilities of AI Scientist systems with human research goals, including but not limited to: (1) Value-driven hypothesis generation: aligning hypothesis proposals with the researcher’s research preferences, risk tolerance and resource constraints. (2) Ethical boundary management: Dynamically adapt to the researcher’s institutional rules, funding ethics, and personal boundaries (e.g., auto-rejecting high-risk topics like gene editing). (3) Collaborative research agendas: Planning research directions with human researchers jointly, taking into account individual career aspirations and priorities for impactful scientific contributions.
@@ 8.2.2 AI Scientist for Human Society
Prospects for human development. AI Scientist systems can greatly boost human progress by: (1) Reducing resource gaps by providing open-access tools and cloud-based experimental simulators, enabling under-resourced institutions to participate in global research. (2) Accelerating discovery for critical challenges (e.g., renewable energy, disease treatment) through automated experimentation. Early successes like ChemCrow (Bran et al., 2023) and BioDiscoveryAgent (Roohani et al., 2024) show promise in chemistry/biology. As these systems increasingly interface with physical laboratories (e.g., IoT sensors), they may evolve to support full-cycle scientific workflows.
@@ 9 Conclusion
This paper presented a systematic survey of existing research on AI Scientist systems, providing a comprehensive overview of advances in the field. Specifically, we proposed a capability-level framework that divides the capabilities of an AI Scientist into four progressive levels: knowledge acquisition, idea generation, verification and falsification, and evolution. Based on this framework, we critically analyzed current research and highlighted gaps AI Scientist systems must address before they can make ground-breaking discoveries that solve grand challenges across medicine, energy, and the environment, thereby changing the world and reshaping the scientific research paradigm. Finally, we concluded the survey by outlining future directions essential for closing the gap and advancing the field toward truly autonomous scientific intelligence.